OpenStudy (anonymous):

METAL/FAN for whoever can help me with these questions FAST i have got to get 8 more assignments in algebra done by tonight and its already 7:51 pm... PLEASE!!!!

jimthompson5910 (jim_thompson5910):

What's your question?

OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = −34 and a9 = −122? −408 −397 −386 −375 What is the 7th term of the geometric sequence where a1 = 256 and a3 = 16? 0.03125 0.0625 0.125 0.15625 Every day, there are 5 times more likes on an internet video of a dog which is modeled by the function c(n) = (5)n − 1, where n is the number of days since the video posted. On the first day, there were 103 likes. What is the function that shows the number of likes each day? c(n) = (5)(103)(n − 1) c(n) = (103)n − 1 c(n) = 103(5)n − 1 c(n) = (5)103 − 1 A new video game is expected to sell 100 copies the first hour at a local game store. After that, the sales will follow the function s(x) = 12(x − 1) where x is the number of hours. What is the function that shows total sales, including the first hour? s(x) = 12(x − 1) − 100 s(x) = 100x − 12 s(x) = 100 − 12(x − 1) s(x) = 100 + 12(x − 1)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok it would be good to post these one at a time so they're easy to reference

jimthompson5910 (jim_thompson5910):

What is the 32nd term of the arithmetic sequence where a1 = −34 and a9 = −122? −408 −397 −386 −375

jimthompson5910 (jim_thompson5910):

what's the common difference in this case?

OpenStudy (anonymous):

When did you receive this assignment?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

now, i am working on it as i type....

OpenStudy (anonymous):

Are you doing online schooling, home school?

OpenStudy (anonymous):

i have like 17 more questions i have left to do.. for this 1 assignment. and yes i am, FLVS

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = -34 + d(n-1) a9 = -34 + d(9-1) -122 = -34 + d(9-1) solve for d to get d = ???

OpenStudy (anonymous):

i think it is 4....

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

okay just give me a sec...

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

okay this time, im pretty sure. is it 6?

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = -34 + d(n-1) a9 = -34 + d(9-1) -122 = -34 + d(9-1) -122 = -34 + d(8) -122 = -34 + 8d -122+34 = 8d -88 = 8d 8d = -88 d = -88/8 d = -11

jimthompson5910 (jim_thompson5910):

Hopefully you'll see where you went wrong

OpenStudy (anonymous):

i honestly was going to say 8 but i doubted myself because i thought no that cant be right! it seemed like too high of a number...

jimthompson5910 (jim_thompson5910):

do you see how I got -11?

OpenStudy (anonymous):

yes, i understand that. (:

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = -34 + d(n-1) an = -34 + (-11)(n-1) an = -34 - 11n + 11 an = -11n - 23

OpenStudy (anonymous):

mhm, so what would be my next step?

jimthompson5910 (jim_thompson5910):

now plug in n = 32

OpenStudy (anonymous):

-375?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

alright now what about the next 3? What is the 7th term of the geometric sequence where a1 = 256 and a3 = 16? 0.03125 0.0625 0.125 0.15625

OpenStudy (anonymous):

that would have to have a big Common difference correct?

jimthompson5910 (jim_thompson5910):

it's not an arithmetic sequence this time

jimthompson5910 (jim_thompson5910):

it's geometric

jimthompson5910 (jim_thompson5910):

so we have a common ratio, call it r

OpenStudy (anonymous):

ahh, see theyre all mixed and it gets me like super overwhelmed.. ive got so much to do...

OpenStudy (anonymous):

alright so how would i do this?

jimthompson5910 (jim_thompson5910):

To jump from term to term, you multiply the last term by r So to go from a1 to a2, you multiply by r a1 = 256 a2 = a1*r a2 = 256*r

OpenStudy (anonymous):

because the answer i got was 17.06666666..

jimthompson5910 (jim_thompson5910):

to go from a2 to a3, you again, multiply by r a3 = a2*r a3 = (256*r)*r a3 = 256*r^2

OpenStudy (anonymous):

well, my problem is how do i find r?

jimthompson5910 (jim_thompson5910):

a3 to a4 a4 = a3*r a4 = (256*r^2)*r a4 = 256*r^3

jimthompson5910 (jim_thompson5910):

see how the pattern is forming?

OpenStudy (anonymous):

yeah, i think i remember doing something like this...

jimthompson5910 (jim_thompson5910):

a3 = 256*r^2 16 = 256*r^2 r = ???

OpenStudy (anonymous):

im not quite sure if my calculator is right it keeps coming up with a decimal..

OpenStudy (anonymous):

and now 1...? ehh

jimthompson5910 (jim_thompson5910):

what solution did you get

OpenStudy (anonymous):

i now got 16? i think that one is correct

jimthompson5910 (jim_thompson5910):

16 = 256*r^2 16/256 = r^2 1/16 = r^2 r^2 = 1/16 r = ???

OpenStudy (anonymous):

i take that as i was wrong... .-.

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so is it 0.0625?

jimthompson5910 (jim_thompson5910):

that's the decimal value of 1/16

jimthompson5910 (jim_thompson5910):

take the square root of that to get???

OpenStudy (anonymous):

omg i didnt realize that the answers in this were decimals.. >.> woow and i think 256?

OpenStudy (anonymous):

alright wanna help me with some more?

jimthompson5910 (jim_thompson5910):

this is what I mean r^2 = 1/16 r = sqrt(1/16) r = 1/4 or r = 0.25

OpenStudy (anonymous):

oh yeah, i have NEVER been good a sqare rooting... EVER no matter how much i try...

jimthompson5910 (jim_thompson5910):

So the nth term is an = a1*r^(n-1) an = 256*(0.25)^(n-1) and then you plug in n = 7 to get the 7th term a7 = 256*(0.25)^(7-1) a7 = 0.0625

jimthompson5910 (jim_thompson5910):

well you have a calculator to help you out

OpenStudy (anonymous):

my calculator has not square root sign...

jimthompson5910 (jim_thompson5910):

what kind of calc do you have?

OpenStudy (anonymous):

its a basic.

jimthompson5910 (jim_thompson5910):

ok so use wolfram alpha as a calculator or you can use google as a calculator

jimthompson5910 (jim_thompson5910):

go ahead, try it type in sqrt(1/16) or any math expression

OpenStudy (anonymous):

i dont have time for that.. maybe after i finish all my assignments i will. i have to get this done by tonight...

jimthompson5910 (jim_thompson5910):

it's as simple as typing "sqrt(1/16)" into google it's not that bad

OpenStudy (anonymous):

ok hold on...

OpenStudy (anonymous):

i got 0.25

jimthompson5910 (jim_thompson5910):

good, so that shows you how r = 1/4 or r = 0.25

OpenStudy (anonymous):

okay, that makes a lot of since.

jimthompson5910 (jim_thompson5910):

I'm glad it does